OpenStudy (konradzuse):

taylor series f(x) = ln(x) c = 1

OpenStudy (konradzuse):

Yeah, that's what I got too... But apparently not correct.....

OpenStudy (experimentx):

let x = u + 1, expand it at u=0

OpenStudy (experimentx):

http://www.math.com/tables/expansion/log.htm

OpenStudy (konradzuse):

well actually the sumation starts at 0.

OpenStudy (konradzuse):

SO maybe ^n-1?

OpenStudy (experimentx):

f(x) = ln (x), x=1 => f(x) = 0 f'(x) = 1/x, x=1 => f(x) = 1 f''(x) = -1/x^2 => f(x) = -1 .. so on and on

OpenStudy (konradzuse):

hmm...

OpenStudy (experimentx):

use the regular formula .. f^n(1)(x-1)^n/n!

OpenStudy (konradzuse):

f^n(1)?

OpenStudy (experimentx):

that's nth derivative of log(x) at x=1

OpenStudy (konradzuse):

Ok, so the same as what wolfram said, except n! instead of n?

OpenStudy (experimentx):

well, you will get some numbers from the derivative that would help cancelling some parts of n!

OpenStudy (experimentx):

it is essentially same as expansion of ln(1+u) at u=0 http://home.scarlet.be/math/taylor.htm#ln(1+x)-and-its-Macl then put back u = x-1

OpenStudy (turingtest):

where are you stuck?

OpenStudy (turingtest):

the Taylor expansion for a function around x=a is\[f(x)=\sum_{n=0}^\infty\frac{f^{(n)}(a)}{n!}(x-a)\]

OpenStudy (turingtest):

in this case a=1 so we have\[\ln x=\sum_{n=0}^\infty\frac{f^{(n)}(1)}{n!}(x-1)\]so for the first term, what is f(1) (that is when n=0) ??